Operations Report – January 11th



Crew 272 Operations Report 11-1-2023

SOL: 10

Name of person filing report: Mason Kuhn

Non-nominal systems: Oven light, Suit 7’s vent panel on the back of the suit (one of the screw holes in the wood)

Notes on non-nominal systems: Noticed the oven light today while cooking dinner. It is definitely not an urgent problem. 1 of 2 screw holes in the wood to secure the vent panel to suit 7 is too large for the washer and nut. Ideally, the wood should be repaired/replaced, but I can temporarily secure the vent panel with a larger diameter washer and nut.


Spirit rover used: No

Hours: 207.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No

Hours: 111.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 219.0

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Perseverance rover used: No

Hours: 253.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes on rovers: None were used on the EVA to Hab Ridge.

Summary of Hab operations: Suit 7 repairs were made on the lower deck, see suits summary for information. 4 showers were taken today, so the water usage was above our average.

WATER USE: 26.1 gallons

Water (static tank): 180.8 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: No

Summary of internet: Nominal

Summary of suits and radios: Suit 7 has been repaired! There was a loose connection in terminal 3 of the toggle switch, and the wire-organizing zip tie was too tight. These items were resulting in spotty functionality. A new zip tie is still organizing the wires inside the suit back, but it is now a bit looser. The terminal 3 connection was retightened to the underside of the toggle switch, and Suit 7 electrical systems are now nominal. Suits 8,9,10,11 were used on the EVA. All suits are charging. All radios have been charged to full and removed from the charging docks.

Summary of GreenHab operations: Monitored and watered GreenHab crops and experimental crops. The door was propped open most of the sol due to warmer temperatures.

WATER USE: 7 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 5g beans

Summary of Science Dome operations: N/A

Dual split: On (from approximately 8:00 pm to 8:00 am)

Summary of RAM operations: Placed the observatory and Science Dome items in their correct locations. Retrieved needle nose pliers, wrench, and screwdriver from the RAM, they will be returned to the RAM immediately.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support:

  • Please advise if you would like me to complete the temporary repair on suit 7’s vent panel, or if you would like me to leave it as-is. The vent panel is secure, but the left side of the panel has about 0.5 inches of undesired movement.

Best Regards,

Mason Kuhn

Crew Engineer

EVA Report – January 11th



Crew 272 EVA Report 11-01-2023

EVA # 8

Author: Arly Black

Purpose of EVA: Exploratory EVA to Hab Ridge, initial scouting of geologic setting.

Start time: 2:00 PM

End time: 4:15 PM

Narrative: Due to unpredictability in the weather and rover charging mishaps, today’s EVA reverted to Plan B: hike to Hab Ridge and explore the area without any preconceptions of what to expect. This time, no rovers were harmed in the making of the EVA. The crew walked from the Hab along road 1103 towards North Ridge. They took a left and walked towards Hab Ridge Road. Along the way, the curious traveler stopped to collect samples and examine their surroundings. They then continued on towards Hab Ridge Road, which runs north and south and lies to the west of the Hab, where they found a ridiculous amount of “devil’s toenails” (a genus of extinct oysters, also known as Gryphaea). They walked until they reached the repeater (it’s still standing!) and were able to see our glorious Hab off in the distance. This was the goal of the mission and therefore mission success was achieved! The crew did note that this EVA required more skillful map reading than any other EVA so far.

As this was an exploratory mission, the crew used the experiences from this EVA to brainstorm ideas for future EVAs (but only two days of EVAs left ☹!!!). Adriana examined the geological environment and tried to see how well defined the strata were in that location and what kinds of fossils were present. She was not expecting to see that many toenails which is indicative of an ancient marine environment in that area (or many giants with nail clippers). She is interested in investigating what they have been preserved in. The strata are not well defined along Hab Ridge Road, and it is difficult to see the contact between formations.

On their way back to the Hab, Adriana did some more digging, as one does, and was excited to find even MORE toenails – who wouldn’t be?! Many photos were taken to document the area and they trundled on back to the Hab along the same route. Well done explorers! But leave those toe clippings off the kitchen table, would ya?

Destination: Hab Ridge

Coordinates (use UTM WGS 84): N517800, E4251500

Participants: Adriana Brown (Crew Geologist), Kshitij Mall (Crew Commander), Madelyn Whitaker (GreenHab Officer), Megan Rush (HSO Officer)

Road(s) and routes per MDRS Map: Walked on foot up Road 1103 until Hab Ridge, then returned on foot by Road 1103 to the Hab.

Mode of travel: Walking

Journalist Report – January 11th



Crew 272 Journalist Report 1-11-2023
Sol: 10
Author: Kenny Pritchard, Crew Journalist
Title: Kissing Hab Ridge

We set off Arly’s CO2/air quality monitor early this morning thanks to breathing exercises led by Kshitij — Anulom, Vilom, and Pranayama, which had the entire crew huffing and puffing (and blowing our noses in between exercises). After the breathing exercises, we engaged in some yoga asans including Shav-asan (corpse-pose), which felt fitting considering the unfortunate quantity of sim-deaths to this point. For breakfast, Madelyn revisited the crew-approved egg casserole recipe from a previous sol. Due to an intense Martian dust storm and some unfavorable ground conditions, one of our two planned EVAs was cancelled today, but four lucky Martians still got to feel fresh air (or at least EVA-suit air) when a new plan was set.

With veteran experience in gearing up for EVA, Adriana, Madelyn, Megan, and Kshitij were out of the Hab in no time, and they followed a road northwest to reach an overlook called Hab Ridge. From this Ridge, they were able to overlook the Hab. Compared to Kissing Camel, this is nomenclature I can get behind! The trip was an exploratory EVA which the crew completed entirely on foot, employing their best map-reading skills to tirelessly traverse the tricky terrain (or masterfully maneuver the merciless ‘marrain?’). Along the way, they happened upon some Gryphaea fossils or ‘devil’s toenails’ – extinct oysters from several hundred million years ago. I reckon they collected enough for a hearty seafood dinner, but nobody seemed to have the appetite for it. From the top of Hab Ridge, the EVA crew had a great view of the Hab (as I mentioned) and the surrounding topography. They got some great pictures from the vantage point and a better navigational understanding of the area, and Adriana will use this practice to plan out the final EVAs for her research.

Once again among the Hab-rats, I took some video documentation of our crew and the Hab to be compiled later. Mason toiled away at fixing the air circulation system in EVA suit #7, to great success! On-site repairs are highly preferable to service shipments back to the manufacturer, which can take years to return. Arly packed out to the Science Dome to clean off petrified wood and other samples she’d gathered, then set to preparing a popular baked-feta pasta recipe from TikTok. With only two full sols remaining before we depart from MDRS, home is on everyone’s minds! Until then, we will make the most of the time we have here.

GreenHab Report – November 25th


Crew 272 GreenHab Report 11-01-2023

GreenHab Officer: Madelyn Whitaker

Environmental control: heater, door propped open

Average temperatures: 74 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 6 gallons

Daily water usage for research and/or other purposes: 1 gallon

Water in Blue Tank: 229 gallons

Time(s) of watering for crops: 1000

Changes to crops: Harvested beans. No new plantings.

Narrative: Monitored all experimental and GreenHab crops. Harvested beans for a snack. Door propped open most of the day due to warmer temperatures.

Harvest: (include which crop and mass in grams): Harvested 5g beans

Support/supplies needed: N/A

Sol Summary Report – January 11th



Hello Mission Support,

Following is the Sol Summary Report.

Crew 272 Sol Summary Report 11-01-2023

Sol: 10

Summary Title: Hab Ridge: The New Frontier

Author’s name: Kshitij Mall

Mission Status: All nominal

Sol Activity Summary: After successfully completing watching the movie Lakshya (Objective), the crew got pumped up to achieve their Lakshya, which is to retrieve BTS possibly displaced by the heavy wind past night. We also planned to traverse the Hab Ridge and identify places of interest for geological experiments for Crew Geologist Adriana. To calm the team down from the sim-death experience last sol with the outer Hab door opening up by a wind blast, I shared Yog comprising some breathing exercises like Pranayama, Anulom, Vilom, some warmup exercises and some asans (poses) with the crew. As per the crew, it did benefit them after being on a streak of EVAs and living in their new isolated Hab Sweet Hab for quite some time. The breakfast for the sol included egg casserole, which was prepared by our Greenhab Officer Madelyn and supported by HSO Megan. A mistake on my part to not plug the chargers back into two of the rovers thinking that overcast conditions will drain the limited solar power we had followed by bad weather in the night made conditions not congenial for carrying out any EVA. But Mission Support recommended us for an exploratory mission to the Hab Ridge, which we were originally planning to be just a geological EVA instead of retrieving my electrical-Martian-baby, BTS. He will have to wait another day for the rescue, but I’m confident he is strong enough to handle more Martian winds for a few sols. The advice from the Mission Support indeed was handy as the sol unfolded. For exploratory missions we didn’t require extensive Earth Mission Support.

After getting an approval from the Mission Support, an exploratory EVA to the Hab Ridge was then conducted at 2 PM by Adriana and me as the primaries for the geology and exploration part, respectively, with Madelyn and Megan as the supporting EVA crew. Good map reading skills were the need of the hour for this EVA. Our EVA crew had great success in reaching the newly established Repeater Station by Crew 271. We also took some great pictures of the EVA crew at the Hab Ridge with Hab Sweet Hab and our rovers in the backdrop. The remaining boilermaker engineers and scientists were busy doing their rock souvenir shopping on Mars while I was taking pictures and videos that could be used by future EVA crews, who plan to visit the Hab Ridge. I too collected some Gryphaea or the Devil’s toenails as Martian omiyage (souvenirs) for my Earth-based families including my 10-month-old son, Hiro.

After coming back from the EVA at 4:15 PM, Adriana and I debriefed the EVA to Arly. In the meantime, Crew Engineer Mason attempted fixing suit #7 that had connection issues and had a great success in fixing suit #7. We also completed the daily human-factors based cognitive performance tests and personal surveys for Arly and Andres. Dinner duties for the sol are handled by Arly and Mason, who are preparing TikTok Pasta. We might play Monikers card game tonight, which we have been procrastinating for several sols. Tomorrow, we plan to conduct a geological EVA to the Cowboy corner to study the paleo flow indictors and do some tool test of 3D printed hammers. Madelyn would be helped tomorrow during her Great Harvest at the Greenhab by the rest of the crew based on their availabilities during the sol.

Look Ahead Plan: Conduct a geological EVA at the Cowboy corner and complete the Great Harvest at the Greenhab

Anomalies in work: None

Weather: Very Sunny, Calm Winds

Crew Physical Status: Good

EVA: One

Reports to be filed: Sol Summary, Operations Report, Journalist Report, EVA Report, EVA Request, Greenhab Report.

Support Requested: None


Copyright © The Mars Society. All rights reserved. | Main Site